OpenStudy (anonymous):

Solving Systems By Substitution y=3x+11 y=-2x+1

Parth (parthkohli):

y equals 3x + 11 y also equals -2x + 11 We can just say \(3x + 11 = -2x + 1\) Solve for x.

OpenStudy (anonymous):

i need alittle more assistance

Parth (parthkohli):

Okay. Aight. \(\Large \color{Black}{\Rightarrow 5x + 11 = 1 }\) Adding 2x to both sides. \(\Large \color{Black}{\Rightarrow 5x = -10 }\) Subtracting both sides by 11. \(\Large \color{Black}{\Rightarrow x = -2}\) Dividing both sides by 5.

Parth (parthkohli):

Now y = 3x + 11 So, y = 3(-2) + 11 y = -6 + 11 = 5